# Alpha-element abundance patterns in star-forming regions of the local Universe

### v1

@@ Appendix D Tables of atomic data, references, physical conditions and abundances
Table 3: References for atomic data used for collisionally excited lines.
Ion Transition probabilities Collision strengths
O+ Froese Fischer & Tachiev (2004) Kisielius et al. (2009)
O2+ Wiese et al. (1996), Storey & Zeippen (2000) Storey et al. (2014)
Ne2+ McLaughlin et al. (2011) McLaughlin et al. (2011)
S+ Irimia & Froese Fischer (2005) Tayal & Zatsarinny (2010)
S2+ Froese Fischer et al. (2006) Grieve et al. (2014)
Cl2+ Fritzsche et al. (1999) Butler & Zeippen (1989)
Ar2+ Mendoza (1983), Kaufman & Sugar (1986) Galavis et al. (1995)
Ar3+ Mendoza & Zeippen (1982) Ramsbottom & Bell (1997)
Fe2+ Deb & Hibbert (2009), Mendoza et al. (2023) Zhang (1996), Mendoza et al. (2023)
Table 4: Identification number, galaxy, region name, references and type of object for the nebulae analysed in this study. The complete table will be available online only.
ID Galaxy Region Reference Type
#1 IC 1613 S3 Bresolin et al. (2007) HII
#2 IC 1613 S7 Bresolin et al. (2007) HII
#3 LMC 30 Doradus Peimbert (2003) HII
#4 LMC IC 2111 Domínguez-Guzmán et al. (2022) HII
#5 LMC N11B Domínguez-Guzmán et al. (2022) HII
#6 LMC N44C Domínguez-Guzmán et al. (2022) HII
#7 LMC NGC 1714 Domínguez-Guzmán et al. (2022) HII
#8 M101 H1013 Bresolin (2007) HII
#9 M101 H1013 Croxall et al. (2016) HII
#10 M101 H1013 Esteban et al. (2009) HII
#11 M101 H1013 Kennicutt et al. (2003) HII
#12 M101 H1018 Croxall et al. (2016) HII
#13 M101 H103 Croxall et al. (2016) HII
#14 M101 H1040 Croxall et al. (2016) HII
#15 M101 H1045 Croxall et al. (2016) HII
#16 M101 H104 Croxall et al. (2016) HII
#17 M101 H1052 Croxall et al. (2016) HII
#18 M101 H1105 Kennicutt et al. (2003) HII
#19 M101 H1118 Esteban et al. (2020) HII
Table 5: Electron densities and temperatures derived for the nebular sample. The complete table will be available online only.
ID Adopted $n_{\rm e}$ $T_{\rm e}$([N ii]) $T_{\rm e}$([S iii]) $T_{\rm e}$([O iii])
[cm-3] [K] [K] [K]
#1 $100\pm 100$ $18270^{+490}_{-660}$
#2 $100\pm 100$ $14290^{+570}_{-530}$
#3 $400^{+20}_{-30}$ $10440^{+220}_{-340}$ $10070^{+80}_{-110}$ $9880\pm 60$
#4 $260\pm 90$ $9790^{+250}_{-260}$ $9330^{+210}_{-250}$ $9100^{+150}_{-130}$
#5 $270^{+90}_{-70}$ $9800^{+130}_{-110}$ $10520^{+240}_{-210}$ $9160^{+110}_{-90}$
#6 $100\pm 100$ $10470^{+330}_{-270}$ $11450^{+320}_{-220}$ $11320^{+190}_{-140}$
#7 $410^{+160}_{-120}$ $10190^{+550}_{-460}$ $9510^{+340}_{-230}$ $9480^{+190}_{-170}$
#8 $100\pm 100$ $8130^{+210}_{-160}$ $7700^{+800}_{-640}$ $7610^{+180}_{-240}$
#9 $100\pm 100$ $7550^{+220}_{-170}$ $6890^{+140}_{-110}$ $7390^{+230}_{-200}$
#10 $100\pm 100$ $7850^{+230}_{-360}$ $7380^{+390}_{-760}$
#11 $100\pm 100$ $7840^{+460}_{-610}$ $6740^{+200}_{-330}$
#12 $100\pm 100$ $7560^{+480}_{-450}$ $7640^{+330}_{-420}$
#13 $100\pm 100$ $9560^{+560}_{-950}$ $11130^{+1280}_{-940}$ $8920^{+300}_{-250}$
#14 $100\pm 100$ $7300^{+460}_{-390}$ $7520^{+230}_{-350}$
#15 $100\pm 100$ $7590^{+430}_{-200}$ $7290^{+320}_{-260}$
#16 $100\pm 100$ $10750^{+360}_{-160}$
#17 $210^{+60}_{-50}$ $8370^{+180}_{-160}$ $8960^{+180}_{-110}$ $7750\pm 70$
#18 $200^{+90}_{-80}$ $9340^{+720}_{-1130}$ $7910^{+260}_{-210}$ $8950^{+190}_{-290}$
#19 $100\pm 100$ $12550^{+450}_{-550}$
Table 6: Ionic abundances derived from the nebular sample. All values are in 12+log(Xi+/H+) units. The complete table will be available online only.
ID O+ O2+ S+ S2+ Ne2+ Ar2+ Ar3+
#1 $6.82^{+0.05}_{-0.04}$ $7.55^{+0.04}_{-0.03}$ $5.12^{+0.14}_{-0.15}$ $7.02^{+0.04}_{-0.04}$ $5.22^{+0.05}_{-0.04}$
#2 $7.60^{+0.07}_{-0.06}$ $7.42^{+0.05}_{-0.04}$ $6.62^{+0.06}_{-0.06}$
#3 $7.60^{+0.05}_{-0.04}$ $8.27^{+0.01}_{-0.01}$ $5.52^{+0.03}_{-0.03}$ $6.63\pm 0.01$ $7.68^{+0.01}_{-0.01}$ $6.02\pm 0.01$ $4.54^{+0.02}_{-0.03}$
#4 $8.06^{+0.05}_{-0.05}$ $8.19\pm 0.03$ $5.83^{+0.04}_{-0.04}$ $6.72^{+0.04}_{-0.04}$ $7.55^{+0.04}_{-0.04}$ $6.02^{+0.04}_{-0.04}$ $4.18\pm 0.08$
#5 $7.98^{+0.03}_{-0.03}$ $8.17\pm 0.02$ $5.80^{+0.02}_{-0.02}$ $6.44^{+0.03}_{-0.03}$ $7.57^{+0.03}_{-0.03}$ $5.87^{+0.03}_{-0.02}$ $4.28\pm 0.02$
#6 $7.31^{+0.06}_{-0.05}$ $8.22^{+0.02}_{-0.02}$ $5.54^{+0.03}_{-0.03}$ $6.25^{+0.03}_{-0.03}$ $7.69\pm 0.03$ $5.76^{+0.03}_{-0.03}$ $5.73^{+0.02}_{-0.02}$
#7 $7.66^{+0.11}_{-0.08}$ $8.28^{+0.04}_{-0.03}$ $5.47^{+0.07}_{-0.05}$ $6.67^{+0.04}_{-0.04}$ $7.71^{+0.05}_{-0.04}$ $6.03^{+0.04}_{-0.04}$ $4.55^{+0.05}_{-0.05}$
#8 $8.33^{+0.06}_{-0.05}$ $8.01^{+0.06}_{-0.05}$ $6.12^{+0.04}_{-0.04}$ $6.85^{+0.16}_{-0.12}$ $7.17^{+0.07}_{-0.07}$ $6.09^{+0.15}_{-0.10}$
#9 $8.45^{+0.07}_{-0.06}$ $8.05^{+0.06}_{-0.05}$ $6.19^{+0.04}_{-0.04}$ $7.12\pm 0.03$ $7.25^{+0.08}_{-0.06}$ $6.30^{+0.03}_{-0.03}$
#10 $8.19^{+0.10}_{-0.08}$ $8.05^{+0.17}_{-0.11}$ $6.11^{+0.06}_{-0.05}$ $6.91^{+0.11}_{-0.09}$ $7.19^{+0.25}_{-0.14}$ $6.14^{+0.06}_{-0.06}$
#11 $8.34^{+0.19}_{-0.12}$ $8.19^{+0.18}_{-0.11}$ $6.16^{+0.10}_{-0.08}$ $7.20^{+0.06}_{-0.05}$ $7.38^{+0.21}_{-0.13}$ $6.32^{+0.07}_{-0.05}$
#12 $8.35^{+0.19}_{-0.11}$ $7.84^{+0.15}_{-0.11}$ $6.46^{+0.10}_{-0.07}$ $6.66^{+0.06}_{-0.06}$
#13 $8.17^{+0.19}_{-0.13}$ $7.89^{+0.06}_{-0.05}$ $6.09^{+0.10}_{-0.08}$ $6.25^{+0.10}_{-0.07}$ $7.08^{+0.07}_{-0.05}$ $5.74^{+0.12}_{-0.08}$
#14 $8.52^{+0.20}_{-0.12}$ $8.13^{+0.17}_{-0.10}$ $6.51^{+0.11}_{-0.07}$ $6.83^{+0.05}_{-0.04}$ $7.38^{+0.19}_{-0.14}$
#15 $8.33^{+0.12}_{-0.08}$ $8.05^{+0.10}_{-0.08}$ $6.26^{+0.07}_{-0.05}$ $6.73^{+0.05}_{-0.05}$ $7.26^{+0.15}_{-0.09}$
#16 $7.97^{+0.05}_{-0.04}$ $7.56^{+0.04}_{-0.03}$ $6.13^{+0.03}_{-0.03}$ $6.29^{+0.03}_{-0.03}$ $6.81^{+0.05}_{-0.04}$ $5.74^{+0.04}_{-0.03}$
#17 $8.14^{+0.05}_{-0.04}$ $8.38^{+0.02}_{-0.02}$ $5.93^{+0.03}_{-0.03}$ $6.65^{+0.03}_{-0.02}$ $7.67^{+0.02}_{-0.02}$ $6.08^{+0.02}_{-0.02}$
#18 $7.97^{+0.26}_{-0.14}$ $8.22^{+0.06}_{-0.04}$ $5.87^{+0.14}_{-0.09}$ $7.00^{+0.04}_{-0.04}$ $7.60^{+0.07}_{-0.05}$ $6.23^{+0.05}_{-0.04}$
#19 $7.66^{+0.08}_{-0.06}$ $7.91^{+0.05}_{-0.04}$ $5.79^{+0.05}_{-0.04}$ $6.50^{+0.16}_{-0.13}$ $7.26^{+0.07}_{-0.06}$ $5.71^{+0.06}_{-0.05}$
Table 7: Total abundances of O and Ne, S and Ar using different ICF schemes. A21: Amayo et al. (2021); I06: Izotov et al. (2006); D13: Dors et al. (2013); D16: Dors et al. (2016) and PM07: Pérez-Montero et al. (2007). All values are in 12+log(X/H) units. The complete table will be available online only.
Ne/H S/H Ar/H
ID O/H A21 I06 D13 A21 I06 D16 A21 I06 PM07
#1 $7.62^{+0.03}_{-0.03}$ $7.12^{+0.04}_{-0.04}$ $7.05^{+0.04}_{-0.04}$ $7.08^{+0.04}_{-0.04}$
#2 $7.82^{+0.04}_{-0.04}$ $7.13^{+0.06}_{-0.06}$ $6.80^{+0.06}_{-0.06}$ $6.84^{+0.06}_{-0.06}$
#3 $8.35^{+0.01}_{-0.01}$ $7.79\pm 0.01$ $7.72^{+0.01}_{-0.01}$ $7.74^{+0.01}_{-0.01}$ $6.72^{+0.01}_{-0.01}$ $6.84\pm 0.01$ $6.75\pm 0.01$ $6.04\pm 0.01$ $6.04\pm 0.01$ $6.05\pm 0.01$
#4 $8.43\pm 0.03$ $7.86^{+0.04}_{-0.03}$ $7.74^{+0.04}_{-0.04}$ $7.69^{+0.04}_{-0.04}$ $6.78^{+0.03}_{-0.03}$ $6.79\pm 0.03$ $6.79^{+0.03}_{-0.03}$ $6.05^{+0.04}_{-0.04}$ $6.05^{+0.04}_{-0.04}$ $6.07^{+0.04}_{-0.04}$
#5 $8.39\pm 0.01$ $7.84^{+0.03}_{-0.03}$ $7.73^{+0.03}_{-0.03}$ $7.69^{+0.03}_{-0.03}$ $6.55^{+0.02}_{-0.02}$ $6.56\pm 0.02$ $6.56^{+0.02}_{-0.02}$ $5.91^{+0.03}_{-0.02}$ $5.90^{+0.02}_{-0.03}$ $5.92^{+0.02}_{-0.02}$
#6 $8.27\pm 0.02$ $7.76\pm 0.03$ $7.69\pm 0.03$ $7.73\pm 0.03$ $6.43^{+0.02}_{-0.02}$ $6.63\pm 0.02$ $6.46\pm 0.02$ $6.05^{+0.02}_{-0.02}$ $6.04\pm 0.02$ $6.06^{+0.02}_{-0.01}$
#7 $8.37^{+0.03}_{-0.03}$ $7.83^{+0.04}_{-0.04}$ $7.76^{+0.05}_{-0.04}$ $7.77^{+0.05}_{-0.04}$ $6.74\pm 0.03$ $6.85\pm 0.04$ $6.77^{+0.04}_{-0.03}$ $6.05^{+0.04}_{-0.03}$ $6.04^{+0.04}_{-0.03}$ $6.06^{+0.03}_{-0.04}$
#8 $8.50^{+0.04}_{-0.04}$ $7.77^{+0.07}_{-0.07}$ $7.56^{+0.07}_{-0.07}$ $7.46^{+0.07}_{-0.07}$ $6.91^{+0.12}_{-0.12}$ $6.90^{+0.12}_{-0.12}$ $6.92^{+0.12}_{-0.12}$ $6.16^{+0.13}_{-0.12}$ $6.16^{+0.15}_{-0.10}$ $6.22^{+0.15}_{-0.10}$
#9 $8.59^{+0.05}_{-0.05}$ $7.87^{+0.07}_{-0.07}$ $7.68^{+0.08}_{-0.06}$ $7.57^{+0.08}_{-0.06}$ $7.17^{+0.03}_{-0.03}$ $7.15\pm 0.03$ $7.17^{+0.02}_{-0.03}$ $6.37^{+0.06}_{-0.05}$ $6.38^{+0.03}_{-0.03}$ $6.45^{+0.03}_{-0.03}$
#10 $8.42^{+0.08}_{-0.07}$ $7.69^{+0.20}_{-0.19}$ $7.49^{+0.25}_{-0.14}$ $7.41^{+0.25}_{-0.14}$ $6.97^{+0.09}_{-0.09}$ $6.96^{+0.09}_{-0.08}$ $6.98^{+0.09}_{-0.08}$ $6.18^{+0.23}_{-0.11}$ $6.19^{+0.06}_{-0.06}$ $6.25^{+0.06}_{-0.06}$
#11 $8.57^{+0.11}_{-0.10}$ $7.87^{+0.16}_{-0.17}$ $7.68^{+0.21}_{-0.13}$ $7.59^{+0.21}_{-0.13}$ $7.24^{+0.14}_{-0.08}$ $7.22\pm 0.05$ $7.24^{+0.05}_{-0.05}$ $6.38^{+0.21}_{-0.11}$ $6.38^{+0.07}_{-0.05}$ $6.43^{+0.07}_{-0.05}$
#12 $8.47^{+0.12}_{-0.12}$ $6.86^{+0.17}_{-0.09}$ $6.85\pm 0.05$ $6.87^{+0.05}_{-0.05}$
#13 $8.35^{+0.10}_{-0.10}$ $7.65^{+0.06}_{-0.06}$ $7.44^{+0.07}_{-0.05}$ $7.34^{+0.07}_{-0.05}$ $6.48^{+0.07}_{-0.06}$ $6.46^{+0.07}_{-0.06}$ $6.48^{+0.07}_{-0.06}$ $5.80^{+0.11}_{-0.09}$ $5.80^{+0.12}_{-0.08}$ $5.87^{+0.12}_{-0.08}$
#14 $8.67^{+0.12}_{-0.12}$ $8.00^{+0.15}_{-0.16}$ $7.80^{+0.19}_{-0.14}$ $7.70^{+0.19}_{-0.14}$ $7.00^{+0.17}_{-0.10}$ $6.98^{+0.04}_{-0.04}$ $7.00^{+0.04}_{-0.04}$
#15 $8.52^{+0.07}_{-0.08}$ $7.83^{+0.11}_{-0.12}$ $7.62^{+0.15}_{-0.09}$ $7.53^{+0.15}_{-0.09}$ $6.85\pm 0.04$ $6.84^{+0.04}_{-0.04}$ $6.86^{+0.04}_{-0.04}$
#16 $8.11^{+0.03}_{-0.04}$ $7.44^{+0.04}_{-0.05}$ $7.19^{+0.05}_{-0.04}$ $7.13^{+0.05}_{-0.04}$ $6.51^{+0.02}_{-0.02}$ $6.50^{+0.02}_{-0.02}$ $6.52^{+0.02}_{-0.02}$ $5.81^{+0.04}_{-0.04}$ $5.81^{+0.04}_{-0.03}$ $5.89^{+0.04}_{-0.03}$
#17 $8.58^{+0.02}_{-0.02}$ $7.92^{+0.02}_{-0.02}$ $7.82^{+0.02}_{-0.02}$ $7.79^{+0.02}_{-0.02}$ $6.74\pm 0.02$ $6.77^{+0.02}_{-0.02}$ $6.75^{+0.02}_{-0.02}$ $6.11^{+0.02}_{-0.02}$ $6.11^{+0.02}_{-0.02}$ $6.14^{+0.02}_{-0.02}$
#18 $8.41^{+0.08}_{-0.08}$ $7.84^{+0.06}_{-0.05}$ $7.75^{+0.07}_{-0.05}$ $7.71^{+0.07}_{-0.05}$ $7.05^{+0.04}_{-0.04}$ $7.07^{+0.04}_{-0.04}$ $7.05^{+0.04}_{-0.04}$ $6.26\pm 0.04$ $6.26^{+0.05}_{-0.04}$ $6.29^{+0.05}_{-0.04}$
#19 $8.10^{+0.04}_{-0.04}$ $7.51^{+0.06}_{-0.07}$ $7.39^{+0.07}_{-0.06}$ $7.38^{+0.07}_{-0.06}$ $6.59\pm 0.12$ $6.62^{+0.12}_{-0.13}$ $6.61^{+0.12}_{-0.12}$ $5.74^{+0.05}_{-0.05}$ $5.74^{+0.06}_{-0.05}$ $5.77^{+0.06}_{-0.05}$

### v2

@@ Appendix D Tables of atomic data, references, physical conditions and abundances
The complete reference tables associated with Appendix D can be found at: https://zenodo.org/records/15005362.
The references for the atomic data used per ion are: O+: Froese Fischer & Tachiev (2004), Kisielius et al. (2009); O2+: Storey & Zeippen (2000), Storey et al. (2014), Wiese et al. (1996); Ne2+: McLaughlin et al. (2011); S+: Irimia & Froese Fischer (2005), Tayal & Zatsarinny (2010); S2+: Froese Fischer et al. (2006), Grieve et al. (2014); Cl2+: Butler & Zeippen (1989), Fritzsche et al. (1999); Ar2+: Galavis et al. (1995), Kaufman & Sugar (1986), Mendoza (1983); Ar3+: Mendoza & Zeippen (1982), Ramsbottom & Bell (1997); Fe2+: Deb & Hibbert (2009), Mendoza et al. (2023), Zhang (1996).
The references for the spectroscopic data are: Arellano-Córdova et al. (2016, 2021); Aver et al. (2022); Berg et al. (2013, 2015, 2016, 2020, 2021); Bresolin (2007, 2011a, 2011b); Bresolin et al. (2004, 2005, 2007, 2009b, 2009a, 2010); Castellanos et al. (2002); Crockett et al. (2006); Croxall et al. (2015, 2016); Delgado-Inglada et al. (2016); Díaz et al. (2000); Domínguez-Guzmán et al. (2022); Edmunds & Pagel (1984); Egorova et al. (2021); Esteban & García-Rojas (2018); Esteban et al. (2004, 2009, 2013, 2014, 2017, 2020); Fernández et al. (2018, 2022); Fernández-Martín et al. (2017); García-Rojas et al. (2004, 2005, 2006, 2007); Garnett & Kennicutt (1994); Garnett et al. (1997); Goddard et al. (2011); Gómez-González et al. (2024); Guseva et al. (2009, 2011, 2012, 2013, 2020, 2024); Hägele et al. (2006, 2008, 2011, 2012); Hsyu et al. (2017, 2018); Isobe et al. (2022); Izotov & Thuan (1998, 2004); Izotov et al. (1997, 1999, 2006, 2009, 2012, 2016, 2017, 2018, 2019, 2020, 2021c, 2021b, 2021a, 2022, 2024b, 2024a); Kehrig et al. (2011); Kennicutt et al. (2003); Kirsanova et al. (2023); Kniazev et al. (2005, 2018); Kojima et al. (2021); Lee et al. (2005); Lin et al. (2017); López-Sánchez et al. (2007, 2015); Luridiana et al. (2002); Magrini et al. (2010); Méndez-Delgado et al. (2021a, b, 2022b); Mesa-Delgado et al. (2009); Nakajima et al. (2022); Pagel et al. (1979, 1980); Patterson et al. (2012); Peimbert (2003); Peimbert et al. (2005, 2012); Peña et al. (2007); Peña-Guerrero et al. (2012); Pustilnik et al. (2005, 2011, 2016, 2020); Rogers et al. (2021, 2022); Rosolowsky & Simon (2008); Skillman (1985); Skillman et al. (2013); Stanghellini et al. (2010); Thuan & Izotov (2005); Thuan et al. (2022); Toribio San Cipriano et al. (2016); Valerdi et al. (2019, 2021); Vilchez et al. (1988b); Watanabe et al. (2024); Zurita & Bresolin (2012).